# Multi-RAG: A Multimodal Retrieval-Augmented Generation System for Adaptive Video Understanding

### v1

@@ II-C Existing Assistants
MerryQurry is an LLM-powered educational assistant designed to provide personalized, trustworthy support for both students and educators. Built on a framework, it delivers course-specific, source-grounded responses while addressing concerns around academic integrity and over reliance on AI [21]. STREAMMIND is a video-language framework designed for real-time streaming video understanding and dialogue. It enables AI systems to continuously watch video streams and proactively generate context-aware, natural language responses without requiring explicit user prompts [8].
@@ III SYSTEM DESIGN
Building on the single primary modality paradigm discussed in Section II, our system adopts this approach to build the video encoder, knowledge database, and RAG pipeline shown in Figure 2. Specifically, it integrates multimodal information from video, audio, and documents, by converting all inputs into unified textual representations [2]. Visual and audio data are processed into descriptive text and auxiliary metadata, which are indexed in a vector database, while the original files are archived for potential future reference. At inference, user queries are matched to relevant content through semantic retrieval, and a large language model generates the final response based on both the query and retrieved information. In the following sections, we detail the video encoder framework, construction of the knowledge database, and the RAG pipeline.
Figure 2: The framework of our Multi-RAG pipeline contains three key modules. In the video encoder, the VLM, receiving input from the frame sampler, and ASR process visual and audio streams, respectively, converting multimodal external inputs into unified textual descriptions and embeddings. Next, these are stored in the knowledge vector database, where a retrieval-augmented generation (RAG) agent selects relevant information in response to user query. The LLM then synthesizes the retrieved content to generate the final system response.
@@ III-A Video Encoder
To determine whether a newly sampled frame contains sufficient visual change for further processing, we compare its MSE against a cutoff value derived from sample videos of a recorded lecture series. In general-purpose scenarios characterized by high-frequency transformations of visual content, the utility derived from input-filtering in the pre-processing stage is notably diminished. Consequently, for tasks of this nature, a uniform sampling methodology is employed. For the given video, the frame sampler samples $n$ frames at a constant rate. These frames $\{F_{0},F_{1},...F_{n}\}$ are fed into Vision Language Model to generate the text descriptions $\{D_{0},D_{1},...D_{n}\}$ with timestamps. In addition to descriptions of the depicted content, auxiliary metadata are also generated, which include summaries and preliminary analyses [17]. Temporal summaries are generated at fixed intervals or upon slide transitions using prompts that ensure preservation of global context and mitigate overemphasis on individual frames. Preliminary analysis explains the video’s principal thematic focus, its predominant stylistic characteristics, and salient scene transitions. These auxiliary outputs augment frame-level descriptions by providing a macroscopic account of content progression alongside an initial interpretive framework of the video’s subject matter and structure.
@@ III-B Knowledge Database
The choice of source file formats and the granularity of retrieval units profoundly influences the quality of subsequent generative outputs. Markdown has emerged as a preferred format for large language models due to several advantages: (1) it imposes a clear hierarchical structure through standardized syntax for headings, lists, and other organizational elements; (2) it supports typographic emphases—such as bolding, italics, and code spans—that facilitate nuanced semantic distinctions; and (3) it remains inherently human-readable, thereby enabling seamless collaboration between automated systems and human reviewers. As such, the text descriptions generated in Section III-A are formatted as Markdown files. To build the knowledge database, the content of these documents is chunked and embedded, or encoded as vectors, for storage within a unified vector space.
@@ III-B2 Embedding
We utilized LangChain to generate embeddings for the contextual documents, enabling seamless integration of the LLM into the retrieval pipeline described in Section III-C [10]. Each text chunk is passed through an embedding model (OpenAIEmbeddings), and the output dense vector captures the chunk’s semantic content in a continuous space. The resulting vectors are stored in a vector database (Chroma).
@@ IV-A Supporting Datasets
The QA setting in MMBench-Video benchmark is structured around a three-level hierarchical capability taxonomy. At the highest level (Level 1) capabilities are divided into two domains: Perception and Reasoning. These are further subdivided into nine Level 2 capabilities: Coarse Perception (CP), Fine-grained Perception with Single-Instance (FP-S), Fine-grained Perception with Cross-Instance (FP-C), Hallucination (HL), Logic Reasoning (LR), Attribute Reasoning (AR), Relation Reasoning (RR), Common Sense Reasoning (CSR), and Temporal Reasoning (TR). Level 3 capabilities consists of 26 fine-grained categories, capturing a broad spectrum of cognitive skills relevant to video understanding.
@@ IV-B Evaluation Procedures
Subsequently, the frame rate was effectively halved to 0.5 fps by removing alternate frames, and a second evaluation was performed. Finally, an Automatic Speech Recognition (ASR) module was employed to transcribe the video’s audio content. This transcribed audio information was then integrated into the respective datasets for both the 1 fps and 0.5 fps conditions. Subsequently, as detailed in Section III-A, documents were augmented with auxiliary contextual information to facilitate a further evaluation of the results. To reduce operational costs and latency, GPT-4.1 mini was chosen over GPT-4o, despite offering no additional improvements in model intelligence. Aligned with the evaluation methodology established in the MMBench-video, we employed GPT-4 to assess the discrepancy between model’s generated answer and ground truth answers.
@@ VI Conclusion
We hope this work fosters deeper interdisciplinary discussions and inspires novel frameworks that push the frontier of continual robot learning—from memory augmentation to human-norm acquisition—ultimately empowering robots to become more capable, collaborative, and socially intelligent partners in our shared future.
@@ Acknowledgments
This work has been partially supported by the U.S. Army DEVCOM Army Research Laboratory Cooperative Agreement No. W911NF2120211.

### v2

@@ II-C Existing Assistants
MerryQurry is an LLM-powered educational assistant that provides personalized, source-grounded support for students and educators, while addressing concerns around academic integrity and AI overreliance [24]. Expanding LLMs to domain-specific tasks, Kallakurik et al. [16] propose a medical AI assistant for edge devices that uses input-driven saliency to prune irrelevant neurons, enabling efficient deployment under hardware constraints. Moving beyond language-only inputs, STREAMMIND introduces a video-language framework for real-time streaming comprehension and dialogue, allowing AI systems to proactively respond to video content using context-aware natural language generation [9].
@@ III SYSTEM DESIGN
Building on the single primary modality paradigm discussed in Section II, our system adopts this approach to build the video encoder, knowledge database, and RAG pipeline shown in Figure 2. Specifically, it integrates multimodal information from video, audio, and documents, by converting all inputs into unified textual representations [2]. Visual and audio data are processed into descriptive text and auxiliary metadata, which are indexed in a vector database, while the original files are archived for potential future reference. At inference, user queries are matched to relevant content through semantic retrieval, and a large language model generates the final response based on both the query and retrieved information. In the following sections, we detail the video encoder framework, the construction of the knowledge database, and the RAG pipeline.
Figure 2: The framework of our Multi-RAG pipeline contains three key modules. In the video encoder, the VLM, receives input from the frame sampler, and the ASR processes visual and audio streams, respectively, converting multimodal external inputs into unified textual descriptions and embeddings. Next, these are stored in the knowledge vector database, where a retrieval-augmented generation (RAG) agent selects relevant information in response to the user query. The LLM then synthesizes the retrieved content to generate the final system response.
@@ III-A Video Encoder
To determine whether a newly sampled frame contains sufficient visual change for further processing, we compare its MSE against a cutoff value derived from sample videos of a recorded lecture series. In general-purpose scenarios characterized by high-frequency transformations of visual content, the utility derived from input-filtering in the pre-processing stage is notably diminished. Consequently, for tasks of this nature, a uniform sampling methodology is employed. For the given video, the frame sampler samples $n$ frames at a constant rate. These frames $\{F_{0},F_{1},...F_{n}\}$ are fed into the Vision Language Model to generate the text descriptions $\{D_{0},D_{1},...D_{n}\}$ with timestamps. In addition to descriptions of the depicted content, auxiliary metadata are also generated, which include summaries and preliminary analyses [19]. Temporal summaries are generated at fixed intervals or upon slide transitions using prompts that ensure preservation of global context and mitigate overemphasis on individual frames. Preliminary analysis explains the video’s principal thematic focus, its predominant stylistic characteristics, and salient scene transitions. These auxiliary outputs augment frame-level descriptions by providing a macroscopic account of content progression alongside an initial interpretive framework of the video’s subject matter and structure.
@@ III-B Knowledge Database
The choice of source file formats and the granularity of retrieval units profoundly influence the quality of subsequent generative outputs. Markdown has emerged as a preferred format for large language models due to several advantages: (1) it imposes a clear hierarchical structure through standardized syntax for headings, lists, and other organizational elements; (2) it supports typographic emphases—such as bolding, italics, and code spans—that facilitate nuanced semantic distinctions; and (3) it remains inherently human-readable, thereby enabling seamless collaboration between automated systems and human reviewers. As such, the text descriptions generated in Section III-A are formatted as Markdown files. To build the knowledge database, the content of these documents is chunked and embedded, or encoded as vectors, for storage within a unified vector space.
@@ III-B2 Embedding
We utilized LangChain[7] to generate embeddings for the contextual documents, enabling seamless integration of the LLM into the retrieval pipeline described in Section III-C [11]. Each text chunk is passed through an embedding model (OpenAIEmbeddings[22]), and the output dense vector captures the chunk’s semantic content in a continuous space. The resulting vectors are stored in a vector database (Chroma[25]).
@@ IV-A Supporting Datasets
The QA setting in the MMBench-Video benchmark is structured around a three-level hierarchical capability taxonomy. At the highest level (Level 1), capabilities are divided into two domains: Perception and Reasoning. These are further subdivided into nine Level 2 capabilities: Coarse Perception (CP), Fine-grained Perception with Single-Instance (FP-S), Fine-grained Perception with Cross-Instance (FP-C), Hallucination (HL), Logic Reasoning (LR), Attribute Reasoning (AR), Relation Reasoning (RR), Common Sense Reasoning (CSR), and Temporal Reasoning (TR). Level 3 capabilities consist of 26 fine-grained categories, capturing a broad spectrum of cognitive skills relevant to video understanding.
@@ IV-B Evaluation Procedures
Subsequently, the frame rate was effectively halved to 0.5 fps by removing alternate frames, and a second evaluation was performed. Finally, an Automatic Speech Recognition (ASR) module was employed to transcribe the video’s audio content. This transcribed audio information was then integrated into the respective datasets for both the 1 fps and 0.5 fps conditions. Subsequently, as detailed in Section III-A, documents were augmented with auxiliary contextual information to facilitate a further evaluation of the results. To reduce operational costs and latency, GPT-4.1 mini was chosen over GPT-4o, despite offering no additional improvements in model intelligence. Aligned with the evaluation methodology established in the MMBench-video, we employed GPT-4 to assess the discrepancy between the model’s generated answer and ground truth answers.
@@ VI Conclusion
We hope this work fosters deeper interdisciplinary discussions and inspires novel frameworks that push the frontier of continual robot learning, from memory augmentation to human-norm acquisition, ultimately empowering robots to become more capable, collaborative, and socially intelligent partners in our shared future.
@@ Acknowledgments
This work has been partially supported by the U.S. Army DEVCOM Army Research Laboratory Cooperative Agreement No. W911NF2120211. We also thank Junsong Lin for providing valuable assistance with the figures in this paper.